# AI in Supply Chain Risk Assessment: A Systematic Literature Review and Bibliometric Analysis

### v2

@@ 7 Challenges and Limitations
    Data Availability and Quality: AI models in SCRA rely heavily on data. However, ensuring the availability and quality of data remains a persistent challenge. Supply chain data often comes from various sources, which can be fragmented, incomplete, or outdated. Researchers must explore ways to consolidate, cleanse, and augment data to enhance the effectiveness of AI models.
@@
    Interpretability and Explainability: Many AI techniques, such as deep learning and neural networks, are often considered as ”black-box” models. This lack of interpretability and explainability is a significant limitation. In SCRA, stakeholders require insights into why a model makes a particular prediction or decision. Developing AI models that are not only accurate but also interpretable is crucial for their adoption in real-world supply chain management.
@@
    Dynamic and Evolving Risk Factors: The supply chain environment is dynamic and continuously evolving. New risk factors emerge, and existing ones change over time. AI models may struggle to adapt to these dynamic conditions. Researchers need to explore techniques for real-time risk assessment and adaptive models that can respond to changing risk profiles.
@@ 8 Future Research Directions
In light of the challenges and limitations, the future research directions in the field of AI-based supply chain risk assessment are multi-faceted and hold the potential to advance the state-of-the-art:
  1. 1.
@@
    AI Model Explainability: Emphasize the development of AI models that deliver accurate predictions and offer comprehensive transparency and interpretability. Techniques like LIME (Local Interpretable Model-Agnostic Explanations) and SHAP (SHapley Additive exPlanations) serve as pivotal approaches to enhance the interpretability of AI models. By employing these methods, the goal is to demystify the decision-making process of AI models, enabling stakeholders to comprehend the rationale behind specific predictions or decisions. This transparency enhances trust and facilitates more effective integration of AI-driven insights into supply chain risk management strategies.
  2. 2.
    Hybrid AI Approaches: Combining multiple AI techniques, such as machine learning and expert systems, can enhance the robustness of supply chain risk assessment. Future research should explore hybrid AI models that leverage the strengths of different approaches to improve prediction accuracy and reliability.
  3. 3.
@@
    Data Quality and Accessibility: Addressing data quality challenges involves multifaceted strategies. Developing robust methods for data cleansing, augmentation, and ensuring high-quality data remains a critical focus. Beyond that, initiatives should be aimed at democratizing access to standardized supply chain data repositories. Creating and maintaining such repositories enables stakeholders across the supply chain spectrum to access reliable and consistent data. This increased accessibility fosters better-informed decision-making processes and promotes using standardized data sets for improved AI model training and validation.
  4. 4.
    Real-time Risk Assessment: Investigating AI models capable of real-time risk assessment will be pivotal. These models should adapt to emerging risks and provide timely alerts or recommendations for risk mitigation.
  5. 5.
    Integration Frameworks: Research on creating integration frameworks for AI-based supply chain risk assessment is essential. These frameworks should facilitate easy integration into existing supply chain management systems, ensuring a smooth transition to AI-driven risk assessment.
  6. 6.
    Cross-Domain Application: Expanding the application of AI in risk assessment to various domains within the supply chain, including logistics, procurement, and production, offers opportunities for comprehensive risk management.
  7. 7.
    Adaptation to Pandemic Dynamics (e.g., COVID-19): Investigate AI models capable of adapting to pandemic-induced disruptions within supply chains. Focus on understanding the unique dynamics and challenges posed by events like COVID-19, aiming to develop AI-driven strategies that enable resilience and agility in such crisis scenarios.
  8. 8.
    Ethical and Legal Considerations: As AI plays an increasingly significant role in decision-making, researchers should also address ethical and legal considerations. This includes issues related to data privacy, bias, and fairness in AI models.
By addressing these challenges and pursuing these research directions, the field of AI-based supply chain risk assessment can make significant strides toward more effective and reliable risk management in an ever-evolving global supply chain landscape.
@@ 9 Conclusions
In this extensive review, we delved into the supply chain risk assessment domain and examined the complex challenges that arise when integrating artificial intelligence, especially machine learning. Our thorough analysis encompassed a substantial pool of 1,717 SCRA papers, ultimately narrowing our focus to a refined subset of 48 papers. This endeavor has yielded significant insights and key findings, making substantial contributions to the SCRA literature.
@@
    Our review thoroughly analyzes the existing state of research concerning the application of AI and ML techniques in SCRM. We have meticulously assessed a substantial body of literature to extract meaningful insights, specifically 48 selected articles.
@@
    Through our analysis, we have identified the AI and ML techniques that are most commonly utilized in the domain of SCRA. This adds clarity to the methodologies employed and underscores the trends shaping the field.
@@
Despite these contributions, our research has some limitations. Our data primarily originated from Google Scholar and Web of Science rather than Scopus, potentially introducing biases. Our focus on papers published between 2014 and 2023 introduces a limitation, as it may omit earlier relevant research that could provide valuable historical context. Additionally, our restriction to English-language papers may have omitted valuable non-English research.
@@
Appendix A Co-authorship Analysis Clusters
Cluster First author
Red cluster(10 items) Gong, Y
Li, I
Ii, Z
Wang, B
Wang, D
Wu, Y
Xu, X
Yang, X
Zhang, H
Zhao, D
Green cluster (7 items) Liu, C
Wang, J
Wang, I
Wang, S
Wang, X
Yang, S
Zhang, X
Blue cluster (5 items) Chen, Y
Ii, J
Liu, S
Liu, Y
Yang, G
Yellow cluster (5 items) Li, X
Lim, MK
Ni, D
Qu, Y
Yang, M
Violet cluster (5 items) Wang, GJ
Xie, C
Yan, XG
Zhou, I
Zhu, Y
Sky-blue cluster (5 items) Hu, X
Wang, Y
Wang, Z
Yao, G
Zhang, Y
Appendix B Co-citation Analysis Clusters
B.1 Cited References
Cluster First author Publication year Journal ISO abbreviation
Red cluster (18items) Baryannis G 2019 Future Gener. Comp.
Baryannis G 2019 Int. J. Prod. Res.
Belhadi A 2021 Ann. Oper. Res.
Breiman I 2001 Machine Learning
Cavalcante IM 2019 Int. J. Inform. Manage.
Chawla NV 2002 J. Artif. Intell. Res.
Choi T.M. 2018 Prod. Oper. Manag.
Garvey MD 2015 Eur. J. Oper. Res.
Ho W 2015 Int. J. Prod. Res.
Klikauer T 2016 Triplec-Commun. Capit.
Sharma R 2020 Comput. Oper. Res.
Tang Cs 2006 Int. J. Prod. Econ.
Tang O 2011 Int. J. Prod. Econ.
Wang G 2016 Int. J. Prod. Econ.
Wichmann P 2020 Int. J. Prod. Res.
Wuttke Da 2013 Int. J. Prod. Econ.
Xu XH 2018 Int. J. Prod. Econ.
Zhu Y 2019 Int. J. Prod. Econ.
Green cluster (13 items) Ambulkar S 2015 J. Oper. Manag.
Barney j. 1991 J. Manage.
Belhadi A 2021 Technol. Forecast. Soc.
Brandon-Jones E 2014 J. Supply Chain Manag.
Chowdhury MMH 2017 Int. J. Prod. Econ.
Christopher M 2004 Int. J. Logist. Manage.
Craighead CW 2007 Desition. Sci.
Dubey R 2021 Int. J. Prod. Res.
Prod El Bazj 2021 Int. J. Prod. Res.
Juttner U 2011 Supply Chain Manag.
Juttner U. 2003 Int. J. Logist-Res. App.
Ponomarov SY 2009 Int. J. Logist. Manag.
Tomlin B 2006 Manage. Sci.
Blue cluster (11items) Chowdhury P 2021 Transport Res. E-Log.
Dwivedi YLC 2021 Int. J. In­Form. Manage.
Ivanov D 2020 Ann. Oper. Res.
Ivanov D 2020 Int. J. Prod. Res.
Ivanov D 2020 Transport Res. E-Log.
Ivanov D 2021 Prod. Plan. Control.
Ivanov Dmitry 2020 Int. J. Integr. Supply Manag.
Min H 2010 Int. J. Logist-Res. App.
Queiroz MM 2020 Ann. Oper. Res.
Toorajipour R 2021 J. Bus. Res.
Van Hoek R 2020 Int. J. Oer. Prod. Man.
Yellow cluster (11 items) Armstrong JS 1977 J. Marketing Res.
Fornell C 1981 J. Marketing. Res.
Giannalcis Mihalis 2016 International Journal Of Production Economics
Grover P 2022 Ann. Oper. Res.
Ivanov D 2019 Int. J. Prod. Res.
Kamble Ss 2020 Int. J. Prod. Econ.
Nguyen T 2018 Comput. Oper. Res.
Papadopoulos T 2017 J. Clean. Prod.
Pettit T 2019 J. Bus. Logist.
Podsa Koff Pm 2003 J. Apple. Psychol.
Saberi S 2019 Int. J. Prod. Res.
Violet cluster (5 items) Dolgui A 2018 Int. J. Prod. Res.
Dolgui A 2019 Int. J. Prod. Res.
Heckmann I 2015 Omega-Int. J. Manag. S.
Hosseini S 2019 Transport Res. E-Log.
Ivanov D 2017 Int. J. Prod. Res.
B.2 Authors
Cluster First author
Red cluster (22 items) Baryannis, G
Brintrup, A
Chopra, S
Christopher, M
Dolgui, A
Fahimnia, B
Ghadge, A
Heckmann, I
Ho, W
Hosseini, S
Ivanov, D
Jabbarzadeh , A
Juttner, U
Kumar, S
Paul, SK
Pettit, TJ
Rajesh, R
Sawik, T
Sheffi, Y
Sodhi, MS
Tang, CS
Wieland, A
Green cluster (19 items) Breiman, I
Chen, J
Chien, D
Choi, TM
Govindan, K
Hofmann, E
Kamble, SS
Kumar, A
Ii, J
Liu, Y
Song, H
Tsao, YE
Tseng, MI
Wang, G
Wang, J
Wang, I
Wang, Y
Zhang, Y
Zhu, Y
Blue cluster (14 items) Akter, S
Bag, S
Belhad, A
Dubey, R
Fomel, C
Gunasekaran, A
Gupta, S
Hair, JF
Min, H
Papadopoulos, T
Podsakoff, PM
Queiroz, MM
Sarkis, J
Wamba, SF
Appendix C Bibliographic Coupling Clusters
C.1 Authors
Cluster First author
Red cluster (15 items) Bag, Surajit
Brintrup, Aleicandra
Choi, Tsan-Ming
Chung, Sai-Ho
Dwivedi, Yogesh
K. Gupta, Shivam
Hosseini, Seyedmohsen
Ivanov, Dmitry
Kalpana, P.
Kar, Arpan Kumar
Kumar, Ajay
Ma, Hoi-Lam
Modgil, Sachin
Raut, Rakesh D.
Wang, Xiaoyi
Green cluster (7 items) Hu, Xiaojian
Wang, Gang-Jin Xie
Xie, Chi
Yan, Xin-Guo
Yao, Gang
Zhou, Li
Zhu,You
Blue cluster (5items) I Xingzhi
Lim, Ming K.
N Du
Qu, Yingchi
Yang, Mei
Yellow cluster (5 items) Chen, Yanju
Liu, Yankui
Liu,Ying
Tsao, Yu-Chung
Yang, Guoqing
Violet cluster (4 items) Chien, Chen-Fu
Hussain, Farookh Khade
Hussain, Omar K.
Saberi, Morteza
Sky-blue cluster (3 items) Bouzembrak, Yamine
Liu, Cheng
Marvin, Hansj P.
C.2 Countries
Cluster Countries
Red cluster (9 items) Austria
Belgium
Germany
Greece
Italy
Lithuania
Netherlands
Russia
Serbia
Green cluster (8 items) Australia
Denmark
Iran
Japan
Peoples Republic of China
Saudi Arabia
Singapore
South Korea
Blue cluster (7 items) Finland
Malaysia
Norway
Pakistan
Scotland
Thailand
United Arab Emirates
Yellow cluster (6 items) Brazil
Canada
India
South Africa
Turkey
USA
Violet cluster (6 items) England
France
Morocco
Poland
Slovenia
Spain
Sky-blue cluster (3items) Qatar
Taiwan
Wales
C.3 Documents
Cluster First author (initial name) (year)
Red cluster (19 items) Bhuiyan (2021)
Chien (2020)
Chung (2021)
Davenport (2020)
Du (2019)
Feuerriegel(2019)
Fu (2019)
Ivanov (2021)
Khan (2019)
Klumpp (2018)
Kumar (2020)
Mortazavi(2015)
Nikolopoulos (2021)
Punia (2020)
Roy (2020)
Salamai (2021)
Sampson (2021)
Villegas (2018)
Wu (2021)
Green cluster (15 items) Behl (2022)
Belhadi
Brintrup (2020)
Cheng (2022)
Duan (2021)
Dubey (2021)
Galetsi (2020)
Gao (2020)
Hopkins (2021)
Ii(2018)
Malik (2022)
Rana (2022)
Raut (2018)
Sobb (2020)
Zouari(2021)
Blue cluster (13 items) Baryannis (2019A)
Cavalcante (2019)
Ebinger (2020)
Hamdi (2018)
Hosseini (2016)
Hosseini (2019)
Hsueh (2015)
Izadikhah (2017)
Jacyna (2020)
Ii (2020)
Mavi(2017)
Svoboda (2021)
Xu (2019)
Violet cluster (12 items) Lin (2019)
Rida (2019)
Rishehchi Fayyaz (2021)
Sang (2021)
Song (2021)
Wang (2020)
Zhang (2021)
Zheng (2021)
Zhu (2016A)
Zhu (2016B)
Zhu (2017)
Zhu (2019)
Sky-blue cluster (12 items) Bag
Bechtsis (2022)
Devarajan
Golan (2021)
Modgil (2022A)
Modgil (2022B)
Moosavi (2022)
Nayal (2022)
Nayal (2023)
Naz (2022)
Queiroz (2022)
Tirkolaee (2021)
Orange cluster (12 items) Afify (2019)
Baryannis (201Gb)
Bouzembrak (201G)
Villegas (2018)
Haghjoo (2020)
Hosseini (2020)
Liu (201G)
Liu (2021)
Ma (2020)
Tordecilla (2021)
Tsao (2020)
Yang (2017)
Brown cluster (8 items) Amjad (2020)
Handfield (2020)
Jesus Saenz (2018)
Kosasih (2022)
Liu (2016)
Protogerou (2021)
Rajesh (2020)
Wichmann (2020)
C.4 Organizations
Clusters Organization (abbreviation)
Red cluster (7 items) Emlyon Business Sch.
Hebei Univ.
Huazhong Univ. Sci. Techn.
Islamic Azad Univ.
Univ. Cambridge
Univ. Technol. Sydney
Univ. Tehran
Green cluster (4 items) Beijing Technol. Business
Chongqing Univ.
Hunan Univ.
Univ. Chinese Acad. Sci.
Blue cluster (3 items) Hong Kong Polytech. Univ.
Indian Inst. Technol. Delhi
Neoma Business Sch.

### v3

@@ 8 Challenges and Limitations
    Data Availability and Quality: AI models in SCRA rely heavily on data. However, ensuring the availability and quality of data remains a persistent challenge. SC data often comes from various sources, which can be fragmented, incomplete, or outdated. Researchers must explore ways to consolidate, cleanse, and augment data to enhance the effectiveness of AI models.
@@
    Interpretability and Explainability: Many AI techniques, such as DL and neural networks, are often considered “black-box” models. This lack of interpretability and explainability is a significant limitation. In SCRA, stakeholders require insights into why a model makes a particular prediction or decision. Developing AI models that are not only accurate but also interpretable is crucial for their adoption in real-world SCM.
@@
    Dynamic and Evolving Risk Factors: The SC environment is dynamic and continuously evolving. New risk factors emerge, and existing ones change over time. AI models may struggle to adapt to these dynamic conditions. Researchers need to explore techniques for real-time risk assessment and adaptive models that can respond to changing risk profiles.
@@ 9 Future Research Directions
In light of the challenges and limitations, future research directions in AI-based SCRA are multi-faceted and have the potential to advance the state-of-the-art. The development of AI models that deliver accurate predictions and offer comprehensive transparency and interoperability should be emphasized. Techniques like Local Interpretable Model-Agnostic Explanations (LIME) and SHapley Additive exPlanations (SHAP) serve as pivotal approaches to enhance the interpretability of AI models. By employing these methods, the goal is to demystify the decision-making process of AI models, enabling stakeholders to comprehend the rationale behind specific predictions or decisions. This transparency enhances trust and facilitates more effective integration of AI-driven insights into SC risk management strategies.
@@
Combining multiple AI techniques, such as ML and expert systems, can enhance the robustness of SCRA. Future research should explore hybrid AI models that leverage the strengths of different approaches to improve prediction accuracy and reliability. Addressing data quality challenges involves multifaceted strategies. Developing robust methods for data cleansing and augmentation and ensuring high-quality data remains a critical focus. Beyond that, initiatives should be aimed at democratizing access to standardized SC data repositories. Creating and maintaining such repositories enables stakeholders across the SC spectrum to access reliable and consistent data. Increased accessibility fosters better-informed decision-making processes and promotes using standardized data sets for improved AI model training and validation. Investigating AI models capable of real-time risk assessment will be pivotal. These models should adapt to emerging risks and provide timely alerts or recommendations for risk mitigation.
Research on creating integration frameworks for AI-based SCRA is essential. These frameworks should facilitate easy integration into existing SC management systems, ensuring a smooth transition to AI-driven risk assessment. Expanding the application of AI in risk assessment to various domains within the SC, including logistics, procurement, and production, offers opportunities for comprehensive risk management. Investigate AI models capable of adapting to pandemic-induced disruptions within SCs. Focus on understanding the unique dynamics and challenges posed by COVID-19, aiming to develop AI-driven strategies that enable resilience and agility in such crisis scenarios. As AI plays an increasingly significant role in decision-making, researchers should also address ethical and legal considerations. This includes issues related to data privacy, bias, and fairness in AI models. By addressing these challenges and pursuing these research directions, AI-based SCRA can make significant strides toward more effective and reliable risk management in an ever-evolving global SC landscape.
@@ 10 Conclusions
This extensive review delved into the SCRA domain and examined the complex challenges of integrating AI, especially ML. Our thorough analysis encompassed a substantial pool of 1,439 SCRA papers, ultimately narrowing our focus to a refined subset of 51 papers. This endeavor has yielded significant insights and key findings, making substantial contributions to the SCRA literature.
@@
    Our review thoroughly analyzes the existing state of research concerning the application of AI and ML techniques in SCRM. We have meticulously assessed a substantial body of literature to extract meaningful insights, specifically 51 selected articles.
@@
    Through our analysis, we have identified the AI and ML techniques that are most commonly utilized in the domain of SCRA. This clarifies the methodologies employed and underscores the trends shaping the field.
@@
Despite these contributions, our research has some limitations. Our data primarily originated from Google Scholar and WoS rather than Scopus, potentially introducing biases. Our focus on papers published between 2015 and 2024 presents a limitation, as it may omit earlier relevant research that could provide valuable historical context. Additionally, our restriction to English-language papers may have omitted valuable non-English research.